# On the rank index of projective curves of almost minimal degree

### v1

@@ 3. Generators of projected curves
Since $H-2F$ is spanned by $\{S^{d_{1}-j-4}T^{j}x\mid 0\leq j\leq d_{1}-4\}\cup\{S^{d_{2}-j-4}T^{j}y\mid 0\leq j\leq d_{2}-4\}$, $I(\pi_{p}(\mathcal{C}_{d}))_{2}/I(\mathcal{S})_{2}$ is generated by quadratic forms on $\mathcal{S}$ corresponding to
@@
[…]
@@
which is a set of quadratic forms whose ranks are at most four. Thus, any quadratic generators of the ideal of $\mathcal{C}$ are generated by quadratic forms of rank at most four, implying that the curve $\pi_{p}(\mathcal{C}_{d})$ satisfies property $\operatorname{QR}(4)$. ∎
@@
Let $e_{c}$ be the $c$-th standard basis for $\mathbb{P}^{d}$ with $c\geq 1$ and $d\geq 4$ Then the corresponding binary form is $f_{e_{c}}=s^{d-c}t^{c}\in\mathbb{K}[s,t]_{d}$ and $f^{\perp}=(S^{d-c+1},T^{c+1})\subset\mathbb{K}[S,T]$. Therefore, by Theorem 3.1, the parametrization of the curve $\mathcal{C}=\pi_{e_{c}}(\mathcal{C}_{d})$ is
@@
$\mathcal{S}$ =$=\{[\alpha^{c-1}\gamma:\alpha^{c-2}\beta\gamma:\cdots:\beta^{c-1}\gamma:\alpha^{d-c-1}\eta:\alpha^{d-c-2}\beta\eta:\cdots:\beta^{d-c-1}\eta]|[\alpha:\beta],[\gamma:\eta]\in\mathbb{P}^{1}\}$
@@
and $\mathcal{C}$ is a section defined by $xT^{c+1}-yS^{d-c+1}$ in $\mathcal{S}$. Since $H-2F$ is spanned by $\{S^{c-j-3}T^{j}x\mid 0\leq j\leq c-3\}\cup\{S^{d-c-j-3}T^{j}y\mid 0\leq j\leq d-c-3\}$, $I(\mathcal{C})_{2}/I(\mathcal{S})_{2}$ is generated by quadratic forms on $\mathcal{S}$ corresponding to
@@
Therefore,
$I(\mathcal{S})=\left(2\times 2\text{ minors of }\begin{bmatrix}z_{0}&z_{1}&\cdots&z_{c-2}&z_{c}&z_{c+1}&\cdots&z_{d-2}\\ z_{1}&z_{2}&\cdots&z_{c-1}&z_{c+1}&z_{c+2}&\cdots&z_{d-1}\end{bmatrix}\right)$
and $I(\mathcal{C})$ is generated by the generators of $I(\mathcal{S})$ and
@@
Since
$I(\pi_{e_{c}}(\mathcal{C}_{d}))=I(\mathcal{C}_{d})\cap\mathbb{K}[z_{0},\dots,z_{c-1},z_{c+1},\dots,z_{d}],$
this gives a constructive proof of the result from [ES96] which states that the elimination of a variable from the ideal of the toric variety $\mathcal{C}_{d}$ results in an ideal generated by binomials.
@@ 4. Monomial projections
In Corollary 3.2, we have seen that the projected curves satisfy property $\operatorname{QR}(4)$. However, we will demonstrate that the projected curves also satisfy property $\operatorname{QR}(3)$ in certain cases.
@@
    If $c=0\text{ or }d$, then $\mathcal{C}_{d,c}$ is a rational normal curve. Therefore, it satisfies property $\operatorname{QR}(3)$ (see [HLMP21]).
@@
    If $c=1\text{ or }d-1$, then $\mathcal{C}_{d,c}$ is a del Pezzo curve. Therefore, it satisfies property $\operatorname{QR}(3)$ (see [Par22]).
Thus, we assume $22\leq c\leq d-$ 2 and $d\geq 6$.
We distinguish between the cases $c=2\text{ or }d-2$ and $33\leq c\leq d-$ 3 because, in the case $c=2\text{ or }d-2$, $\mathcal{C}_{d,c}$ is not generated by quadrics, whereas for $33\leq c\leq d-$ 3, the curve $\mathcal{C}_{d,c}$ is generated by quadrics. For $\mathcal{C}_{d,2}$ (or $\mathcal{C}_{d,d-2}$), we instead consider the scheme $X_{d}=\mathcal{C}_{d,2}\cup\mathbb{L}$ where $\mathbb{L}$ is the trisecant line to $\mathcal{C}_{d,2}$. Note that $X_{d}$ is the scheme generated by quadrics on $\mathcal{C}_{d,2}$.
As we aim to determine whether the curves $\mathcal{C}=\pi_{p}(\mathcal{C}_{d})\subset\mathbb{P}^{d-1}$ satisfy property $\operatorname{QR}(3)$ and to provide additional numerical information beyond simply determining if the property holds, we introduce the following notation:
@@
There is a combinatorial way to find a generating set of $I(\mathcal{C}_{d,c})$. Suppose
$I(\mathcal{C}_{d})=\left(2\times 2\text{ minors of }\begin{bmatrix}x_{0}&x_{1}&\cdots&x_{d-1}\\ x_{1}&x_{2}&\cdots&x_{d}\end{bmatrix}\right)\subset\mathbb{K}[x_{0},x_{1},\dots,x_{d}],$
and define the index of a quadratic monomial $x_{i}x_{j}$ by $i+j$. Observe that
@@
Therefore, the differences between any pairs of quadratic monomials $(x_{i}x_{j},x_{k}x_{l})$ with same index $i+j=k+l$ for $i,j,k,l\neq c$ are generators of $I(\mathcal{C}_{d,c})_{2}$. Moreover, for each integer $0\leq t\leq 2d$, we can obtain #{quadratic monomials whose index is $\#\{\text{quadratic monomials whose index is }t\}-1$ linearly independent quadratic forms for each index.
Among the list of monomials, there are $d+1$ quadratic monomials whose support contains $x_{c}$. For $11\leq c\leq d-$ 1, we lose exactly $d+1$ quadratic generators when eliminating $x_{c}$ from $I(\mathcal{C}_{d})$. Therefore, we have $\binom{d}{2}-(d+1)=\binom{d-1}{2}-2$ linearly independent quadratic generators in $I(\mathcal{C}_{d,c})_{2}$. These generators span $I(\mathcal{C}_{d,c})_{2}$ because $\dim_{\mathbb{K}}I(\mathcal{C}_{d,c})_{2}=\binom{d-2}{2}+d-4=\binom{d-1}{2}-2$ where $I(\mathcal{C}_{d,c})_{2}=I(\mathcal{S})_{2}+(I(\mathcal{C}_{d,c})_{2}/I(\mathcal{S})_{2})$.
Let $X$ be a projective scheme defined by quadratic forms and $I$ be the saturated ideal defining $X$.
@@
    For quadratic forms $q$ and $q^{\prime}$ in $I$, we write
    $q\sim_{I}q^{\prime}$
    if $q-q^{\prime}$ can be expressed as a sum of rank 3 generators of $I$. In particular,
@@
      if $q\sim_{I}0$, then $q$ is spanned by rank 3 quadratic forms in $I$.
@@
    For $0\leq c\leq d$, let $I_{d,c}$ denote the ideal of $\mathcal{C}_{d,c}=\pi_{e_{c}}(\mathcal{C}_{d})$. We write
    $q\sim_{(d,c)}q^{\prime}$
    for quadratic forms $q$ and $q^{\prime}$ in $I_{d,c}$ if $q-q^{\prime}$ can be expressed as a sum of rank 3 generators in $(I_{d,c})_{2}$.
  3. (3)
    $\delta(X,t)$ denotes the number of independent quadratic generators of $I$ whose rank is at most $t$.
We first claim that $\mathcal{C}_{d,c}$ for $d\geq 6$ and $33\leq c\leq d-$ 3 satisfies property $\operatorname{QR}(3)$. To show this claim, we begin by proving a series of computational lemmas.
@@
Let $i$, $j$, $k$, and $l$ be integers such that $0\leq i<j\leq d$, $i<j-k$, $i+j+k\equiv 0\pmod{2}$, and $c\neq i,i+k,j-k,j,\frac{i+j-k}{2},\frac{i+j+k}{2}$. Then,
@@
We fix the decomposition of the line bundle $\mathcal{O}_{\mathbb{P}^{1}}(d)=A^{\otimes 2}\times B$ where
@@
Then
@@
For each $60\leq i\leq d-$ 6, the two quadratic forms $x_{i}x_{i+5}-x_{i+1}x_{i+4}$ and $x_{i+1}x_{i+6}-x_{i+2}x_{i+5}$ satisfy
$(x_{i}x_{i+5}-x_{i+1}x_{i+4})\sim_{(d,i+3)}0\text{ and }(x_{i+1}x_{i+6}-x_{i+2}x_{i+5})\sim_{(d,i+3)}0.$
@@
For a decomposition $\mathcal{O}_{\mathbb{P}^{1}}(d)=\mathcal{O}_{\mathbb{P}^{1}}(2)^{\otimes 2}\otimes\mathcal{O}_{\mathbb{P}^{1}}(d-4)$ of the line bundle,
@@
for $a\in\mathbb{K}$. Since $x_{i}x_{i+6}+x_{i+2}x_{i+4}-2x_{i+1}x_{i+5}\sim_{(d,i+3)}(x_{i}+x_{i+4})(x_{i+2}+x_{i+6})-(x_{i+1}+x_{i+5})^{2}$ and the forms $x_{i}x_{i+2}-x_{i+1}^{2}$, $x_{i}x_{i+4}-x_{i+2}^{2}$, $x_{i+2}x_{i+6}-x_{i+4}^{2}$, and $(x_{i}+x_{i+4})(x_{i+2}+x_{i+6})-(x_{i+1}+x_{i+5})^{2}$ are rank 3 quadratic generators in $I_{d,i+3}$, it follows that
$02a^{3}(x_{i}x_{i+5}-x_{i+1}x_{i+4})-4a^{5}(x_{i+1}x_{i+6}-x_{i+2}x_{i+5})\sim_{(d,i+3)}$ 0
@@
For each $50\leq i\leq d-$ 5, the two quadratic forms $x_{i}x_{i+5}-x_{i+2}x_{i+3}$ and $x_{i+2}x_{i+5}-x_{i+3}x_{i+4}$ satisfy
@@
For a decomposition $\mathcal{O}_{\mathbb{P}^{1}}(d)=\mathcal{O}_{\mathbb{P}^{1}}(2)^{\otimes 2}\otimes\mathcal{O}_{\mathbb{P}^{1}}(d-4)$ of the line bundle,
@@
For $d\geq 6$ and $33\leq c\leq d-$ 3, the ideal of $\mathcal{C}_{d,c}$ is generated by quadratic forms and satisfies property $\operatorname{QR}(3)$.
@@
Table 1. List of monomials given indices.
The index $2,4,8,10$ generators are $x_{0}x_{2}-x_{1}^{2},x_{0}x_{4}-x_{2}^{2},x_{2}x_{6}-x_{4}^{2},x_{4}x_{6}-x_{5}^{2}$ and they are already rank $3$ quadratic generators. The remaining generators are index $5,6,7$ generators. By Lemma 4.3 with $i=0$, index $5$ generator $x_{0}x_{5}-x_{1}x_{4}\sim_{(6,3)}0$ and index $7$ generator $x_{1}x_{6}-x_{2}x_{5}\sim_{(6,3)}0$. By Lemma 4.2, $x_{0}x_{6}+x_{2}x_{4}-2x_{2}x_{4}=x_{0}x_{6}-x_{2}x_{4}\sim_{(6,3)}0$ and $x_{0}x_{6}+x_{2}x_{4}-2x_{1}x_{5}=(x_{0}x_{6}-x_{2}x_{4})+2(x_{2}x_{4}-x_{1}x_{5})\sim_{(6,3)}0$. Hence $I_{6,3}$ satisfies $\operatorname{QR}(3)$ property.
@@
The remaining monomials are $x_{0}x_{7}$, $x_{1}x_{7}$, and $x_{2}x_{7}$. We need to find quadratic generators containing these monomials that can be expressed as a sum of rank $3$ quadrics without using $x_{3}$. By Lemma 4.2, we have $x_{0}x_{7}+x_{2}x_{5}-2x_{1}x_{6}\sim_{(7,3)}0$. The generator $x_{1}x_{7}-x_{4}^{2}$ is already a rank $3$ quadratic form. In addition, we obtain $x_{2}x_{7}-x_{4}x_{5}\sim_{(7,3)}0$ by applying Lemma 4.4 with $i=2$. Thus, we conclude that $I_{7,3}$ satisfies property $\operatorname{QR}(3)$.
@@
The remaining monomials are $x_{0}x_{8}$, $x_{1}x_{8}$, and $x_{2}x_{8}$. The binomials $x_{0}x_{8}-x_{4}^{2}$ and $x_{2}x_{8}-x_{5}^{2}$ are already rank $3$ generators. By Lemma 4.2, we have $x_{1}x_{8}+x_{2}x_{7}-2x_{4}x_{5}\sim_{(8,3)}0$. Thus, $I_{8,3}$ satisfies property $\operatorname{QR}(3)$.
@@
The remaining monomials are $x_{0}x_{8}$, $x_{1}x_{8}$, $x_{2}x_{8}$, and $x_{3}x_{8}$. By Lemma 4.2, we have $x_{0}x_{8}+x_{2}x_{6}-2x_{3}x_{5}\sim_{(8,4)}0$ and $x_{1}x_{8}+x_{3}x_{6}-2x_{2}x_{7}\sim_{(8,4)}0$. In addition, the binomial $x_{2}x_{8}-x_{5}^{2}$ is already a rank $3$ quadratic form. Moreover, we obtain $x_{3}x_{8}-x_{5}x_{6}\sim_{(8,4)}0$ by Lemma 4.4 with $i=3$. Hence, $I_{8,4}$ satisfies property $\operatorname{QR}(3)$.
For the inductive hypothesis, assume that the ideal $I_{d-1,c}$ satisfies property $\operatorname{QR}(3)$ for $d\geq 9$ and $c$ such that $3\leq c\leq\lfloor\frac{d}{2}\rfloor$. Since $I_{d,c}$ contains the generators of $I_{d-1,c}$, we only need to find the rank $3$ quadratic forms that involve $x_{d}$. The generators of index $\geq d+c$ are those of a rational normal curve, so the remaining task is to check the quadratic generators with indices between $d$ and $d+c-1$. We will consider the cases based on the parity of $d$ and $c$.
@@
$d+c-2$$\left\{\begin{array}[]{lr}x_{c-2}x_{d}+x_{c-1}x_{d-1}-2x_{\frac{d+c-3}{2}}x_{\frac{d+c-1}{2}},&\text{for even }c\\ x_{c-2}x_{d}-x_{\frac{d+c-2}{2}}^{2},&\text{for odd }c\\ \end{array}\right.$
@@
Since $\frac{d-1}{2}>c$, all of binomals in Table 2 can be expressed as sums of rank 3 quadrics by Lemma 4.2.
If $d+c-1$ is an even number, the quadratic form $x_{c-1}x_{d}-x_{\frac{d+c-1}{2}}^{2}$ is already a rank 3 quadratic form. If $d+c-1$ is odd, we use the following binomial:
@@
The relation holds because $c<d-4$ implied by $c\leq\frac{d-1}{2}$ and $d\geq 9$.
Case 1-2 : $d$ is odd and $c=\frac{d-1}{2}$. The only issue arises at index $d$ because $x_{\frac{d-1}{2}}=x_{c}$. To resolve this, we use the following binomial:
@@
$d+c-2$$\left\{\begin{array}[]{lr}x_{c-2}x_{d}-x_{\frac{d+c-2}{2}}^{2},&\text{for even}c\\ x_{c-2}x_{d}+x_{c-1}x_{d-1}-2x_{\frac{d+c-3}{2}}x_{\frac{d+c-1}{2}},&\text{forodd }c\\ \end{array}\right.$
@@
Since $\frac{d}{2}>c$, all binomials listed in Table 3 are spanned by rank 3 quadratic forms by Lemma 4.2. For the index $d+c-1$, if $c$ is odd, the form $x_{c-1}x_{d}-x_{\frac{d+c-1}{2}}^{2}$ is already a rank $3$ quadratic generator. However, if $c$ is even, then we can use
@@
The issue arises at indices $d$ and $d+1$ because $x_{\frac{d}{2}}=x_{c}$. To resolve this, we replace these forms with the following binomials:
@@
Proof of Theorem 1.1.
Combine Corollary 3.2 and Theorem 4.5. ∎
Now, we prove that $X_{d}=\mathcal{C}_{d,2}\cup\mathbb{L}$ does not satisfy property $\operatorname{QR}(3)$ for $d\geq 5$.
@@
Let $X_{d}$ be the projective scheme defined by the quadrics on $\mathcal{C}_{d,2}$ for $d\geq 5$. Then, we have the following:

### v2

@@ 3. Generators of projected curves
Since H2FH-2F is spanned by $\{S^{d_{1}-j-4}T^{j}x\mid 0\leq j\leq d_{1}-4\}\cup\{S^{d_{2}-j-4}T^{j}y\mid 0\leq j\leq d_{2}-4\}$, $I(\pi_{p}(\mathcal{C}_{d}))_{2}/I(\mathcal{S})_{2}$ is generated by quadratic forms on $\mathcal{S}$ corresponding to
@@
$\cup\;$ $\{S^{d_{2}-j-4}T^{j}g_{1}(S,T)xy-S^{d_{2}-j-4}T^{j}g_{2}(S,T)y^{2}\mid 0\leq j\leq d_{2}-4\}.$
Here, if $d_{1}\leq 3$ (resp. $d_{2}\leq 3$), then the first (resp. second) set is empty. Since we assume $d_{1}+d_{2}\geq 6$ and $d_{1}\leq d_{2}$, the only possible case with $d_{1}\leq 3$ or $d_{2}\leq 3$ is $(d_{1},d_{2})=(3,3)$, in which both sets are empty.
@@
which is a set of quadratic forms whose ranks are at most four. Thus, any quadratic generator of the ideal of $\mathcal{C}$ is a sum of quadratic forms of rank at most $4$. This implies that the curve $\pi_{p}(\mathcal{C}_{d})$ satisfies property $\operatorname{QR}(4)$. ∎
@@
Let $e_{c}$ be the c-th standard basis for $\mathbb{P}^{d}$ with $d\geq 4$ and $1\leq c\leq\lfloor\frac{d}{2}\rfloor$. Then the corresponding binary form is given by $f:=f_{e_{c}}=s^{d-c}t^{c}\in\mathbb{K}[s,t]_{d}.$ Therefore, by Theorem˜3.1, the parametrization of the curve $\mathcal{C}=\pi_{e_{c}}(\mathcal{C}_{d})$ is
@@
$\mathcal{S}$ $=\{[\alpha^{c-1}\gamma:\alpha^{c-2}\beta\gamma:\cdots:\beta^{c-1}\gamma:\alpha^{d-c-1}\eta:\alpha^{d-c-2}\beta\eta:\cdots:\beta^{d-c-1}\eta]|[\alpha:\beta],[\gamma:\eta]\in\mathbb{P}^{1}\}$
@@
and $\mathcal{C}$ is a section defined by $xT^{c+1}-yS^{d-c+1}$ in $\mathcal{S}$. Since H2FH-2F is spanned by $\{S^{c-j-3}T^{j}x\mid 0\leq j\leq c-3\}\cup\{S^{d-c-j-3}T^{j}y\mid 0\leq j\leq d-c-3\}$, $I(\mathcal{C})_{2}/I(\mathcal{S})_{2}$ is generated by quadratic forms on $\mathcal{S}$ corresponding to
@@
and $I(\mathcal{C})$ is generated by $I(\mathcal{S})$ together with
@@
In fact, since
$I(\mathcal{C})=I(\mathcal{C}_{d})\cap\mathbb{K}[y_{0},\dots,y_{c-1},y_{c+1},\dots,y_{d}],$
where
[…]
this gives a constructive proof of a special case of the result in [5], which states that eliminating a variable from the ideal of a toric variety yields an ideal generated by binomials.
@@ 4. Monomial projections
In Corollary˜3.2, we have seen that the projected curves satisfy property $\operatorname{QR}(4)$. However, we will demonstrate that the projected curves also satisfy property $\operatorname{QR}(3)$ in certain cases.
@@
    If $c=0\text{ or }d$, then $\mathcal{C}_{d,c}$ is a rational normal curve. Therefore, it satisfies property $\operatorname{QR}(3)$ (see [8]).
@@
    If $c=1\text{ or }d-1$, then $\mathcal{C}_{d,c}$ is a del Pezzo curve. Therefore, it satisfies property $\operatorname{QR}(3)$ (see [16]).
Thus, we assume $2\leq c\leq d-2$ and $d\geq 6$.
We distinguish between the cases $c=2\text{ or }c=d-2$ and $3\leq c\leq d-3$ because, in the case $c=2\text{ or }d-2$, $\mathcal{C}_{d,c}$ is not generated by quadrics, whereas for $3\leq c\leq d-3$, the curve $\mathcal{C}_{d,c}$ is generated by quadrics. For $\mathcal{C}_{d,2}$ (or $\mathcal{C}_{d,d-2}$), we instead consider the scheme $X_{d}=\mathcal{C}_{d,2}\cup\mathbb{L}$ where $\mathbb{L}$ is the trisecant line to $\mathcal{C}_{d,2}$. Note that $X_{d}$ is the scheme generated by quadrics on $\mathcal{C}_{d,2}$.
As we aim to determine whether the curves $\mathcal{C}_{d,c}=\pi_{c}(\mathcal{C}_{d})\subset\mathbb{P}^{d-1}$ satisfy property $\operatorname{QR}(3)$ and to provide additional numerical information beyond simply determining whether the property holds, we introduce the following notation:
@@
  1. (A)
    There is a combinatorial way to find a generating set of $I(\mathcal{C}_{d,c})$. Suppose I(C_d) = ( 2 ×2 minors of [x0x1⋯xd-1x1x2⋯xd] ) ⊂K[x_0,x_1,…,x_d], and define the index of a quadratic monomial $x_{i}x_{j}$ by $i+j$. Observe that
@@
    Therefore, the differences between all pairs of quadratic monomials $(x_{i}x_{j},x_{k}x_{l})$ with same index $i+j=k+l$ for $i,j,k,l\neq c$ are generators of $I(\mathcal{C}_{d,c})_{2}$. Moreover, for each integer $0\leq t\leq 2d$, let $u$ be the number of quadratic monomials $x_{i}x_{j}$ of index $t=i+j$ with $i,j\in\{0,1,\dots,d\}\setminus\{c\}$. Then one obtains u1u-1 linearly independent generators of $I(\mathcal{C}_{d,c})_{2}$.
    Among all polynomials there are $d+1$ quadratic monomials whose support contains $x_{c}$. For $1\leq c\leq d-1$, we lose exactly $d+1$ quadratic generators when eliminating $x_{c}$ from $I(\mathcal{C}_{d})$. Therefore, we have $\binom{d}{2}-(d+1)=\binom{d-1}{2}-2$ linearly independent quadratic generators in $I(\mathcal{C}_{d,c})_{2}$. These generators span $I(\mathcal{C}_{d,c})_{2}$ because $\dim_{\mathbb{K}}(I(\mathcal{C}_{d,c}))=\binom{d-2}{2}+d-4=\binom{d-1}{2}-2=\tfrac{1}{2}(d^{2}-3d-2).$
  2. (B)
    Let $X\subseteq\mathbb{P}^{d}$ be a projective scheme defined by quadratic forms and $I$ be the saturated ideal defining $X$.
@@
      For quadratic forms $q$ and $q^{\prime}$ in $I$, we write q ∼_I q’ if $q-q^{\prime}$ can be expressed as a sum of rank 3 generators of $I$. In particular,
@@
        If $q\sim_{I}0$, then $q$ is a sum of rank $3$ quadratic forms in $I$.
@@
      For $0\leq c\leq d$, let $I_{d,c}$ denote the ideal of $\mathcal{C}_{d,c}=\pi_{e_{c}}(\mathcal{C}_{d})$. We write q ∼_(d,c) q’ for quadratic forms $q$ and $q^{\prime}$ in $I_{d,c}$ if $q-q^{\prime}$ can be expressed as a sum of rank 3 generators in $(I_{d,c})_{2}$.
We first claim that $\mathcal{C}_{d,c}$ for $d\geq 6$ and $3\leq c\leq d-3$ satisfies property $\operatorname{QR}(3)$. To show this claim, we begin by proving a series of computational lemmas. We recall from Section˜2.4, given a decomposition of the line bundle $\mathcal{O}_{X}(1)=A^{2}\otimes B$ of $X$,
$Q_{A,B}:H^{0}(X,A)\times H^{0}(X,A)\times H^{0}(X,B)\longrightarrow I(X)_{2}$
is defined by
$Q_{A,B}(s,t,h)=\varphi(s\otimes s\otimes h)\varphi(t\otimes t\otimes h)-\varphi(s\otimes t\otimes h)^{2}.$
@@
Let $i$, $j$, $k$, and $l$ be integers such that $0\leq i<j\leq d$, i<jki<j-k, $i+j+k\equiv 0\pmod{2}$, and $c\neq i,i+k,j-k,j,\frac{i+j-k}{2},\frac{i+j+k}{2}$. Then,
@@
We fix the decomposition of the line bundle $\mathcal{O}_{\mathbb{P}^{1}}(d)=A^{2}\otimes B$ where
@@
Then, in the notation introduced at the section 2.4 we have
@@
For each $0\leq i\leq d-6$, the two quadratic forms $x_{i}x_{i+5}-x_{i+1}x_{i+4}$ and $x_{i+1}x_{i+6}-x_{i+2}x_{i+5}$ satisfy
$(x_{i}x_{i+5}-x_{i+1}x_{i+4})\sim_{(d,i+3)}0\quad\text{and}\quad(x_{i+1}x_{i+6}-x_{i+2}x_{i+5})\sim_{(d,i+3)}0.$
@@
For a decomposition $\mathcal{O}_{\mathbb{P}^{1}}(d)=\mathcal{O}_{\mathbb{P}^{1}}(2)^{2}\otimes\mathcal{O}_{\mathbb{P}^{1}}(d-4)$ we have (in the notation introduced at the section 2.4)
@@
for $a\in\mathbb{K}$. Since $x_{i}x_{i+6}+x_{i+2}x_{i+4}-2x_{i+1}x_{i+5}\sim_{(d,i+3)}(x_{i}+x_{i+4})(x_{i+2}+x_{i+6})-(x_{i+1}+x_{i+5})^{2}$ and since the forms $x_{i}x_{i+2}-x_{i+1}^{2}$, $x_{i}x_{i+4}-x_{i+2}^{2}$, $x_{i+2}x_{i+6}-x_{i+4}^{2}$ and $(x_{i}+x_{i+4})(x_{i+2}+x_{i+6})-(x_{i+1}+x_{i+5})^{2}$ are rank three quadratic generators of $I_{d,i+3}$, it follows that
$2a^{3}(x_{i}x_{i+5}-x_{i+1}x_{i+4})-4a^{5}(x_{i+1}x_{i+6}-x_{i+2}x_{i+5})\sim_{(d,i+3)}0$
@@
For each $0\leq i\leq d-5$, the two quadratic forms $x_{i}x_{i+5}-x_{i+2}x_{i+3}$ and $x_{i+2}x_{i+5}-x_{i+3}x_{i+4}$ satisfy
@@
For a decomposition $\mathcal{O}_{\mathbb{P}^{1}}(d)=\mathcal{O}_{\mathbb{P}^{1}}(2)^{2}\otimes\mathcal{O}_{\mathbb{P}^{1}}(d-4)$ we have (in the notation introduced at the end of Section˜2.4),
@@
For $d\geq 6$ and $3\leq c\leq d-3$, the ideal of $\mathcal{C}_{d,c}$ is generated by quadratic forms and satisfies property $\operatorname{QR}(3)$.
@@ 99 $x_{4}x_{5}$
Table 1. List of monomials with given indices.
The index $2,4,8,10$ generators are $x_{0}x_{2}-x_{1}^{2},x_{0}x_{4}-x_{2}^{2},x_{2}x_{6}-x_{4}^{2},x_{4}x_{6}-x_{5}^{2}$ and they are already rank $3$ quadratic generators. The remaining generators are index $5,6,7$ generators. (See Table˜1.) By Lemma˜4.3 with $i=0$, index $5$ generator $x_{0}x_{5}-x_{1}x_{4}\sim_{(6,3)}0$ and index $7$ generator $x_{1}x_{6}-x_{2}x_{5}\sim_{(6,3)}0$. By Lemma˜4.2, $x_{0}x_{6}+x_{2}x_{4}-2x_{2}x_{4}=x_{0}x_{6}-x_{2}x_{4}\sim_{(6,3)}0$ and $x_{0}x_{6}+x_{2}x_{4}-2x_{1}x_{5}=(x_{0}x_{6}-x_{2}x_{4})+2(x_{2}x_{4}-x_{1}x_{5})\sim_{(6,3)}0$. Hence $I_{6,3}$ satisfies property $\operatorname{QR}(3)$.
@@
The remaining monomials are $x_{0}x_{7}$, $x_{1}x_{7}$, and $x_{2}x_{7}$. We need to find quadratic generators containing these monomials that can be expressed as a sum of rank $3$ quadrics without using $x_{3}$. By Lemma˜4.2, we have $x_{0}x_{7}+x_{2}x_{5}-2x_{1}x_{6}\sim_{(7,3)}0$. The generator $x_{1}x_{7}-x_{4}^{2}$ is already a rank $3$ quadratic form. In addition, we obtain $x_{2}x_{7}-x_{4}x_{5}\sim_{(7,3)}0$ by applying Lemma˜4.4 with $i=2$. Thus, we conclude that $I_{7,3}$ satisfies property $\operatorname{QR}(3)$.
@@
The remaining monomials are $x_{0}x_{8}$, $x_{1}x_{8}$, and $x_{2}x_{8}$. The binomials $x_{0}x_{8}-x_{4}^{2}$ and $x_{2}x_{8}-x_{5}^{2}$ are already rank $3$ generators. By Lemma˜4.2, we have $x_{1}x_{8}+x_{2}x_{7}-2x_{4}x_{5}\sim_{(8,3)}0$. Thus, $I_{8,3}$ satisfies property $\operatorname{QR}(3)$.
@@
The remaining monomials are $x_{0}x_{8}$, $x_{1}x_{8}$, $x_{2}x_{8}$, and $x_{3}x_{8}$. By Lemma˜4.2, we have $x_{0}x_{8}+x_{2}x_{6}-2x_{3}x_{5}\sim_{(8,4)}0$ and $x_{1}x_{8}+x_{3}x_{6}-2x_{2}x_{7}\sim_{(8,4)}0$. In addition, the binomial $x_{2}x_{8}-x_{5}^{2}$ is already a rank $3$ quadratic form. Moreover, we obtain $x_{3}x_{8}-x_{5}x_{6}\sim_{(8,4)}0$ by Lemma˜4.4 with $i=3$. Hence, $I_{8,4}$ satisfies property $\operatorname{QR}(3)$.
For the inductive hypothesis, assume that the ideal $I_{d-1,c}$ satisfies property $\operatorname{QR}(3)$ for $d\geq 9$ and $c$ such that $3\leq c\leq\lfloor\frac{d}{2}\rfloor$. Since $I_{d,c}$ contains the generators of $I_{d-1,c}$, we only need to find the rank $3$ quadratic forms that involve $x_{d}$. The generators of index $\geq d+c$ are those of a rational normal curve, so the remaining task is to check the quadratic generators with indices between $d$ and d+c1d+c-1. We will consider the cases based on the parity of $d$ and $c$.
@@
[…]
@@
Since $\frac{d-1}{2}>c$, all binomials in Table˜2 can be expressed as sums of rank 3 quadrics by Lemma˜4.2.
If d+c1d+c-1 is an even number, the quadratic form $x_{c-1}x_{d}-x_{\frac{d+c-1}{2}}^{2}$ is already a rank 3 quadratic form. If d+c1d+c-1 is odd, we use the following binomial:
@@
The relation holds because c<d4c<d-4, as implied by $c\leq\frac{d-1}{2}$ and $d\geq 9$.
Case 1-2 : $d$ is odd and $c=\frac{d-1}{2}$.
Since $c=\frac{d-1}{2}$, all quadrics in Table˜2 except the index $d$ part does not contain any $x_{c}$. So they can be expressed as sums of rank $3$ quadratic generators by Lemma˜4.2. The only issue arises at index $d$. To resolve this, we replace the form with the following quadratic generator without $x_{c}$ from Lemma˜4.2:
@@
[…]
@@
Since $\frac{d}{2}>c$, all quadrics listed in Table˜3 are spanned by rank 3 quadratic forms by Lemma˜4.2. For the index d+c1d+c-1, if $c$ is odd, the form $x_{c-1}x_{d}-x_{\frac{d+c-1}{2}}^{2}$ is already a rank $3$ quadratic generator. However, if $c$ is even, then we can use
@@
Since $c=\frac{d}{2}$, all quadrics in Table˜3 except index $d$ and $d+1$ does not contain any $x_{c}$. So they can be expressed as sums of rank $3$ quadratic generators by Lemma˜4.2. The issue arises only at indices $d$ and $d+1$. To resolve this, we replace these forms with the following quadratic generators from Lemma˜4.2:
@@
Proof of Theorem˜1.1.
Combine Corollary˜3.2 and Theorem˜4.5. ∎
Now, we consider the case $c=2$. In this case, we have $\operatorname{rank}_{\mathcal{C}_{d}}(p)=3$ and hence $\mathcal{C}$ admits a unique 33–secant line $\mathbb{L}$ and the scheme $X_{d}:=\mathbb{L}\cup\mathcal{C}$ is defined by the quadrics contained in the vanishing ideal $I(\mathcal{C})$ the curve $\mathcal{C}$ (see the comment prior to Theorem 1.3). Our aim is to show, that in this case the scheme $X_{d}$ does not satisfy property $\operatorname{QR}(3)$.
@@
Let $X_{d}$ be the projective scheme defined by the quadrics on $\mathcal{C}_{d,2}$ for $d\geq 5$. Then we have the following: